How much must you deposit each year into your retirement account starting now and continuing through year 11 if you want to be able to withdraw $70,000 per year forever, beginning 31 years from now? Assume the account earns interest at 9% per year.

The amount to be deposited is determined to be $
Answer #1

Money withdrawn beginning 31 years from now = 70000/0.09 = 777777.777

Future value at the beginning of 10th year = 777777.77*(P/F,9%,19) = 777777.77*(0.19449) = 151270

The amount to be deposited = 151270*(A/F,9%,12) = 151270*(0.049651) = 7510.70